OpenStudy (joachim):

Evaluate the line integral 7xy^4ds, where C is the right half of the circle x^2 + y^2 = 9.

OpenStudy (joachim):

What are the steps I need to take to solve this problem?

OpenStudy (irishboy123):

Find ds Do you know how?

OpenStudy (turingtest):

I'm a bit rusty on line integrals, but it seems to be that it should be simple if we convert it to polar coordinates

OpenStudy (joachim):

I have started parametrizing x, y \[x = 3\cos(t)\] and \[y = 3\sin(t)\]

OpenStudy (turingtest):

seems legit. what are the bounds on t? what is ds?

OpenStudy (joachim):

I think t is bounded as follows: \[\frac{ -\pi }{ 2 } \le t \le \frac{ \pi }{ 2 }\] is that correct?

OpenStudy (turingtest):

I agree

OpenStudy (joachim):

I am unsure how I need to continue

OpenStudy (turingtest):

as IrishBoy said, you need ds This is basically the arc length formula: do you know it?

OpenStudy (joachim):

ah of course, so I just calculate the arclenth and the integral is something like this? \[17 \int\limits_{-\pi/2}^{\pi/2} 3\cos(t)(3\sin(t))^43dt\]

OpenStudy (turingtest):

yep :)

OpenStudy (joachim):

cool thanks for your help :)

OpenStudy (turingtest):

my pleasure!

OpenStudy (joachim):

constant should by 7